# Conceptualization, Methodology, Writing, Visualization

### v1

@@ 2.1.1 Voltage
Based on the characteristics of the power system, voltage is closely related to reactive power, which is more difficult to generate and inject into grids than active power sayed2022electric . Thus, the high demand for reactive power from EV nonlinear charging can significantly impact the real-time voltage of grid operation. The direct impact of large-scale EV charging on distribution system operation is a voltage drop at the coupling point, which may lead to voltage deviation exceeding the regulatory requirements. For instance, rahman2022comprehensive investigates the factors affecting voltage distribution, such as locations of power sources and EV penetration levels, and compares conditions of multiple parallel load lines with unequal loads. Furthermore, continuous violations may lead to grid operation instability and could even result in blackouts. Therefore, when voltage violations or frequency violations exceed specified limits, it’s necessary to adopt corrective measures to restore violations to normal levels in order to avoid damage to power equipment and negative effects on grid operation safety.
Regarding voltage regulation, reactive power dispatch and load demand management are effective methods to control voltage drops wu2012pev ; leemput2015reactive ; sousa2015multi . Some studies wang2012load ; li2012impacts compare the effects of random charging and smart charging on the voltage variation of the distribution system, finding that smart charging can increase EV penetration rates. Moreover, another impact on voltage is the issue of three-phase voltage unbalance, which is primarily caused by single-phase charging. Uneven distribution of charging loads across three phases may lead to a severe condition of three-phase voltage unbalance yong2015review . richardson2010impact performs a detailed study on this issue by connecting all EVs to a single phase and confirms the seriousness of the problem. To address this problem, a proper load management strategy is usually adopted to mitigate three-phase voltage unbalance by evenly dispatching EV charging loads across three phases. Figure 2 presents the process of voltage deviation from EV charging and regulating voltage to the normal level kersting2018distribution . In Figure 2, $V_{s}$, $V_{load}$ and $V_{manage}$ represent the initial voltage of the node, the voltage after applying EV charging and loads, and the voltage after exerting management methods, respectively. $RI$ and $jXI$ represent the real and imaginary parts of the impedance of EV charging and loads, respectively. $RI_{cap}$ and $jXI_{cap}$ represent the real and imaginary parts of the impedance of the reactive power compensation supply, respectively.
Figure 2: The phasor diagram for regulating voltage deviation to the normal level.
@@ 2.1.2 Harmonics
Power electronic devices, essential for EV charging, are extensively utilized in this field. However, their use can lead to power quality issues in the distribution system, such as the introduction of harmonics, which affect real-time operations.
As the foundation of EV charging, power electronic devices have been widely used. This has caused power quality degradation in the real-time operation of the distribution system, such as harmonics yong2015review . nguyen2013harmonic shows similar outcomes: EV charging injects significant harmonics into the distribution system and can cause unacceptable Total Harmonic Distortion (THD) of voltage. Regarding studies on the effect of harmonics, establishing models of EV chargers is a critical point, where gatta2016pq develops a sensitivity analysis about the composition of harmonic disturbances due to AC/DC converters installed in EVs, and rahman2022comprehensive compares simulation results for different power converter topologies in steady-state operation.
The basic configuration of an EV charger uses a back-to-back converter structure. On the input side, a diode bridge rectifier determines the current, which is highly peaked and dominated by low-order harmonics. This leads to voltage deviations, reduced power quality, and de-rating of system components yong2015review ; rahman2022comprehensive . The peak current is superimposed on the sinusoidal current drawn by the EV charger and other loads in the distribution system. It produces a non-sinusoidal voltage drop across the grid impedance. Thus, both the coupling point and the distribution grid contain additional harmonics. The effect of the above condition depends on the parameters of the distribution line; if the grid impedance is small, the voltage drop at the coupling point due to non-sinusoidal current is small. Although EV charging may bring harmonic pollution to the power grid, employing filtering and advanced power electronic devices can alleviate this problem. For instance, rai2019bridgeless proposes a single-ended primary inductor converter for power factor correction operation, and rahman2022comprehensive adopts a boost power factor correction circuit along with the diode bridge rectifier to solve the above condition, through improving and regulating the rectified voltage to generate a minimal ripple DC voltage.
@@ 2.2.1 Peak loads
Both electricity consumption and EV charging loads are closely related to human activities. Without controlled charging, EV charging demand may overlap with existing peak loads in the grid fachrizal2020smart . This overlap would further intensify grid loads during peak time, thereby stressing grid operation and posing risks to the security and stability of the distribution system. Peak load increase generated by EV charging has become a critical factor for grid operation and risk assessment stiasny2021sensitivity . Some studies have focused on this issue. For instance, weiller2011plug suggests that, during commuting hours, new load peaks could exceed natural peaks if EV charging loads are not sufficiently shifted to off-peak periods. Another study shafiq2020reliability indicates that uncontrolled EV charging, especially during peak time, could lead to up to 6.89% load loss. Since traditional distribution systems are designed to handle peak loads, reducing peak demand can also significantly lower overall construction costs. Key factors affecting peak loads in the distribution system include EV charging time, charging location, charging power, and penetration rates of EV charging. Thus, the load management strategy can help balance power loads and reduce the difference between peak and valley loads fachrizal2020smart . Common load management methods include off-peak and valley-filling charging, which shifts EV charging from peak to lower-demand times. This charging approach avoids charging during peak periods and fills low consumption periods, reducing system loss and improving load factor sortomme2010coordinated .
@@
Regarding the impact of EV-grid-connected charging on peak loads in practical applications, many countries, including the United States and Germany, have analyzed the effects of EV charging on load distribution based on their specific circumstances and have proposed corresponding solutions. These solutions include delayed charging methods weiller2011plug , using EVs as stable power storage devices in the grid hartmann2011impact , shifting EV charging to nighttime hours mullan2011modelling , transferring EV loads from peak to off-peak periods by implementing demand response (DR) strategies park2013impact and utilizing V2G reverse power flow to reduce peak load in the grid. Specifically, mosaddegh2017optimal developed a Distribution Optimal Power Flow (DOPF) model incorporating a neural network model of controllable loads to mitigate peak loads. Based on a microgrid energy management system framework, solanki2015including proposes optimal dispatch strategies for dispatchable generators, energy storage systems, and controllable peak loads to achieve peak load dispatch effectively.
@@ 2.2.2 The coping capacity of EV charging areas to grid faults
Faults in grids may cause fast variations of voltage and bursts of harmonics, as well as a serious influence on EV charging. For instance, faults causing undervoltage at or below 0.3 p.u. can lead to ceasing charging for 2–10 seconds mckillop2023impact . onar2010grid compares conditions of a three phase-ground fault and after fault clearance, obtaining the conclusion that systems connected to EVs are more sensitive to disturbances and less stable in magnitude deviations and adjustment time. mckillop2023impact investigates the impact of EVs on power grids during network faults through testing EVs’ responses to a double-line-to-ground transmission fault, obtaining the result of successive over-frequency and under-frequency of 50.78 and 49.22 Hz, respectively. Therefore, regarding fault conditions in grids, the ability of EV charging areas to ride through grid faults and mitigate grid faults is worth attention. EV charging areas with low-voltage ride-through (LVRT) function can effectively handle faults and prevent system instability onar2010grid , and V2G mode can also be used for grid support during faults according to LVRT requirements katic2019impact . For instance, katic2019impact tests six characteristic types of faults to demonstrate the positive impact of LVRT. Table 1 summarizes the research on enhancing EV charging areas’ coping capacity to grid faults.
@@
Ref. no. Year Technical foundation Implementation approach Aims and resolved issues
mishra2022intelligent 2022 Least mean square algorithm A reconfigurable multi-objective charging control architecture Overcome various grid abnormalities during EV charging operation
baghaee2019anti 2020 Support vector machines Anti-islanding protection scheme for low voltage-sourced converter-based microgrids Islanding and grid-fault detection.
shin2012development 2012 Wireless sensor network Smart grid monitoring system connected with EV charging system using anti-islanding method EV charging process continues without any serious fault
onar2010grid 2018 Dynamic combination of EV chargers and single-phase induction motors Implement an LVRT scheme to inject reactive power into the grid and regulate EV charging rate Handle faults and prevent dynamic voltage instability.
almeida2015electric 2015 Combination of inertial emulation and droop control Primary frequency control technique with EVs Safe integration of intermittent renewable energy sources
pandit2024frequency 2024 MM-SFR model Propose a non-linear optimization framework incorporating constraints of EV aggregator, frequency security, converter voltage, security, and LVRT constraints Develop a framework for quantifying EVs’ contribution to providing frequency support
mckillop2023impact 2023 Fault condition test Test EVs physically under various network fault conditions using a grid simulator supply interface Characterize fault ride-through (FRT) performances
katic2019impact 2019 V2G mode according to the LVRT requirements Test the possibility of voltage improvement during voltage dips in V2G mode Provide grid support during faults through V2G mode
yadav2023low 2023 LVRT Applications of LVRT Boost the resilience of the power network against extreme events
falahi2013potential 2013 V2G services Pair up a photovoltaic source with an EV charger through a single-phase bidirectional charger topology Study the potential of EVs to help PV sources during LVRT
dietmannsberger2017simultaneous 2017 A three-phase inverter model Inverters with both LVRT capability and anti-islanding protection simultaneously Solve the conflict between LVRT capability and anti-islanding detection requirements
@@ 3.1 Smart charging
In response to the growing number of EVs and their impacts on infrastructure, it is necessary to implement intelligent and coordinated charging management methods engel2022hierarchical . Many studies have focused on optimizing charging management through smart charging approaches, typically integrated with an energy management system (EMS), which is designed to determine optimal charging schemes and regulation strategies that lead to positive effects on the distribution network. Figure 3 presents the negative impacts from EV charging on net loads, and the positive impacts from smart charging. Through establishing models for controllable loads solanki2015including ; hafez2016integrating ; mosaddegh2017optimal , these models are integrated into the distribution system operation framework or EMS solanki2017sustainable ; wi2013electric ; van2016energy ; solanki2015including to determine the optimal charging strategies and dispatch decisions wi2013electric ; van2016energy ; engel2022hierarchical ; solanki2015including ; hafez2016integrating ; kiviluoma2011methodology ; zheng2018online . The advantage of this approach is optimizing EV charging management together with other components of power system. Determining the optimal charging strategies and optimal dispatch decisions is typically formulated as an optimization problem considering multiple factors, with the charging management objective encoded as a cost function engel2022hierarchical . Specifically, in solanki2015including , an EMS framework is proposed to determine optimal scheduling decisions considering dispatchable generators, energy storage systems, and peak demand for controllable loads. zheng2018online transforms the optimal charging problem to an optimal power flow problem to minimize the total system energy cost, then utilizes a modified convex relaxation technique to obtain the globally optimal solution. The key issue in establishing EV load models is reducing the uncertainties caused by EV charging behaviors. Due to the complexity of controllable loads and the limited data, it is difficult to model these loads using basic physical laws mosaddegh2017optimal . Common methods include probabilistic approaches papadopoulos2010predicting ; ramadhani2020review , stochastic optimization van2016energy ; liao2015dispatch , and evolutionary algorithms, such as neural networks hafez2016integrating ; mosaddegh2017optimal , particle swarm optimization regulski2014estimation , or genetic algorithms ju1996nonlinear . Moreover, based on these technical foundations, many studies have focused on developing efficient and accurate real-time management models to address the impact of EV integration on the real-time operation of distribution networks. A detailed introduction to this topic is provided in Section 4.
@@
Figure 3: (a) The effects of EV charging on the electricity load profile on a day in residential distribution grids. (b) The positive impacts of managing EV charging on net loads through PV coordination and smart charging methods. Figure are developed with data from berg2022data .
@@ 3.2 Charging environment management
In addition to charging stations, common EV charging environments include workplaces and households, where rooftop photovoltaic (PV) systems are typically integrated with energy storage systems (ESS). Regarding workplace conditions, the net-zero energy building (NZEB) is a popular topic, and the synergy between EV loads and PV generation in buildings can help achieve this goal fachrizal2020smart . In the context of building energy management system with EV charging stations, engel2022hierarchical proposes a hierarchical economic model predictive control scheme for EV charging management, considering the objectives of monetary costs, building temperature comfort, EV charge satisfaction, and battery degradation. van2016energy designs a workplace energy management system with photovoltaic generation prediction and power flow optimization between PV systems, grids, and battery electric vehicles. lakshminarayanan2018real proposes a real-time optimal EM controller for V2G integration to provide an optimal schedule for the operation of the workplace microgrid system.
The issue of EV charging in households is closely related to the home energy management system (HEMS), which is necessary for residential electricity consumers to participate in DR programs actively ghazvini2017demand . shao2012grid proposes a DR strategy, and ghazvini2017demand presents an intelligent HEMS algorithm to schedule controllable appliances, i.e., EVs in smart homes. In addition, wi2013electric considers user preferences for determining EV charging schedules in smart homes, and shao2012grid introduces consumer comfort indices to measure the impact of demand response on consumers’ lifestyles.

### v2

@@ Abstract
{highlights}
Large-scale integration of electric vehicles introduces negative impacts on the distribution system’s real-time operations.
By leveraging the bidirectional flow of information and energy in smart grids, the adverse effects of EV charging can be minimized and even converted into beneficial outcomes through effective real-time management strategies.
In-depth analysis of the real-time management system for EV charging is conducted by focusing on real-time state estimation of the distribution networks and the management of EV charging activities.
@@ 2.1.1 Voltage
Based on the characteristics of the power system, voltage is closely related to reactive power, which is more difficult to generate and inject into grids than active power sayed2022electric . EV charging is a typical non-linear load based on rectifier circuits and power converters. Thus, the high demand for reactive power from non-linear EV charging can significantly affect the real-time voltage profile of grid operation, which reduces the power factor and exacerbates voltage distortion and fluctuations. Transmitting large amounts of reactive power from generators to loads results in significant transmission losses in the power lines. The direct impact of large-scale EV integration is a voltage drop at the coupling point, which may lead to voltage deviation exceeding the regulatory requirements.
For instance, rahman2022comprehensive investigates the factors affecting voltage distribution, such as locations of power sources and EV penetration levels, and compares conditions of multiple parallel load lines with unequal loads. Furthermore, continuous violations may lead to grid operation instability and could even result in blackouts. Therefore, when voltage violations or frequency violations exceed specified limits, it’s necessary to adopt corrective measures to restore violations to normal levels in order to avoid damage to power equipment and negative effects on grid operation safety.
Regarding voltage regulation, reactive power dispatch and load demand management are effective methods to control voltage drops wu2012pev ; leemput2015reactive ; sousa2015multi . Some studies wang2012load ; li2012impacts compare the effects of random charging and smart charging on the voltage variation of the distribution system, finding that smart charging can increase EV penetration rates. Moreover, another impact on voltage is the issue of three-phase voltage unbalance, which is primarily caused by single-phase charging. Uneven distribution of charging loads across three phases may lead to a severe condition of three-phase voltage unbalance yong2015review . richardson2010impact performs a detailed study on this issue by connecting all EVs to a single phase and confirms the seriousness of the problem. To address this problem, a proper load management strategy is usually adopted to mitigate three-phase voltage unbalance by evenly dispatching EV charging loads across three phases.
@@ 2.1.2 Harmonics
Power electronic devices are extensively utilized in EV charging and discharging. However, their use can lead to power quality issues, such as harmonics, which affects real-time operation of distribution systems yong2015review . nguyen2013harmonic shows similar outcomes: EV charging injects significant harmonics into the distribution system and can cause unacceptable total harmonic distortion of voltage. Regarding studies on the effect of harmonics, establishing models of EV chargers is a critical point, where gatta2016pq develops a sensitivity analysis about the composition of harmonic disturbances due to AC/DC converters installed in EVs, and rahman2022comprehensive compares simulation results for different power converter topologies in steady-state operation.
The basic configuration of an EV charger uses a back-to-back converter structure. On the input side, a diode bridge rectifier determines the current, which is highly peaked and dominated by low-order harmonics. This leads to voltage deviations, reduced power quality, and de-rating of system components yong2015review ; rahman2022comprehensive . The peak current is superimposed on the sinusoidal current drawn by the EV charger and other loads in the distribution system. It produces a non-sinusoidal voltage drop across the grid impedance. Thus, both the coupling point and the distribution grid contain additional harmonics. The effect of the above condition depends on the parameters of the distribution line. If the grid impedance is small, the voltage drop at the coupling point due to non-sinusoidal current is small. Although EV charging may bring harmonic pollution to the power grid, employing filtering and advanced power electronic devices can alleviate this problem. For instance, rai2019bridgeless proposes a single-ended primary inductor converter for power factor correction operation, and rahman2022comprehensive adopts a boost power factor correction circuit along with the diode bridge rectifier to solve the above condition, through improving and regulating the rectified voltage to generate a minimal ripple DC voltage.
@@ 2.2.1 Peak loads
Both electricity consumption and EV charging loads are closely related to human activities wang2022cyber . Without controlled charging, EV charging demand may overlap with existing peak loads in the grid fachrizal2020smart . This overlap would further intensify grid loads during peak time, thereby stressing grid operation and posing risks to the security and stability of the distribution system. Peak load increase generated by EV charging has become a critical factor for grid operation and risk assessment stiasny2021sensitivity . Some studies have focused on this issue. For instance, weiller2011plug suggests that, during commuting hours, new load peaks could exceed natural peaks if EV charging loads are not sufficiently shifted to off-peak periods. Another study shafiq2020reliability indicates that uncontrolled EV charging, especially during peak time, could lead to up to 6.89% load loss. Since traditional distribution systems are designed to handle peak loads, reducing peak demand can also significantly lower overall construction costs. Key factors affecting peak loads in the distribution system include EV charging time, charging location, charging power, and penetration rates of EV charging. Thus, the load management strategy can help balance power loads and reduce the difference between peak and valley loads fachrizal2020smart . Common load management methods include off-peak and valley-filling charging, which shifts EV charging from peak to lower-demand times. This charging approach avoids charging during peak periods and fills low consumption periods, reducing system loss and improving load factor sortomme2010coordinated .
@@
Regarding the impact of EV-grid-connected charging on peak loads in practical applications, many countries, including the United States and Germany, have analyzed the effects of EV charging on load distribution based on their specific circumstances and have proposed corresponding solutions. These solutions include delayed charging methods weiller2011plug , using EVs as stable power storage devices in the grid hartmann2011impact , shifting EV charging to nighttime hours mullan2011modelling , transferring EV loads from peak to off-peak periods by implementing demand response (DR) strategies park2013impact and utilizing V2G reverse power flow to reduce peak load in the grid hu2019prediction . Specifically, shao2012grid establishes a real-time energy management optimization model for an EV parking lot based on a peak load limitation oriented DR program to maximize the load factor. The simulation results in the distribution circuit in Blacksburg show that the proposed DR strategy can maintain the original peak demand with different EV penetration levels. In binetti2015scalable , a scalable real-time greedy algorithm is used to coordinated charging strategies, which reduces the peak value to 10709 MW, compared with the base profile peak of 16327 MW. hu2019prediction simulates V2G mode to reduce peak load on a real low voltage network in England, where the power curve is levelled off at 20% penetration and the maximum penetration level is 50%. mosaddegh2017optimal developed a distribution optimal power flow model incorporating a neural network model of controllable loads to mitigate peak loads. Based on a microgrid energy management system framework, solanki2015including proposes optimal dispatch strategies for dispatchable generators, energy storage systems, and controllable peak loads to achieve peak load dispatch effectively.
@@ 2.2.2 The coping capacity of EV charging areas to grid faults
Faults in grids may cause fast variations of voltage and bursts of harmonics, as well as a serious influence on EV charging. For instance, a fault causing undervoltage at or below 0.3 p.u. can cease charging for 2–10 seconds mckillop2023impact . onar2010grid compares conditions of a three phase-ground fault and after fault clearance, obtaining the conclusion that systems connected to EVs are more sensitive to disturbances and less stable in magnitude deviations and adjustment time. mckillop2023impact investigates the impact of EVs during network faults through testing EVs’ responses to a double-line-to-ground transmission fault, obtaining the result of successive over-frequency and under-frequency of 50.78 and 49.22 Hz, respectively. Therefore, regarding fault conditions in grids, the ability of EV charging areas to ride through grid faults and mitigate grid faults is worth attention. EV charging areas with low-voltage ride-through (LVRT) function can effectively handle faults and prevent system instability onar2010grid , and V2G mode can also be used for grid support during faults according to LVRT requirements katic2019impact . For instance, katic2019impact tests six characteristic types of faults to demonstrate the positive impacts of LVRT. Table 1 summarizes the research on enhancing EV charging areas’ coping capacity to grid faults.
@@
Ref. no. Year Technical foundation Implementation approach Aims and resolved issues Performance metrics
mishra2022intelligent 2022 Least mean square algorithm A reconfigurable multi-objective charging control architecture Overcome various grid abnormalities during EV charging operation The LVRT operation is shown during 0.7-0.8s
baghaee2019anti 2020 Support vector machines Anti-islanding protection scheme for low voltage-sourced converter-based microgrids Islanding and grid-fault detection. Islanding detection is achieved within 45-60 msec
shin2012development 2012 Wireless sensor network Smart grid monitoring system connected with EV charging system using anti-islanding method EV charging process continues without any serious fault The operation of micro-grid system is performed well
onar2010grid 2018 Dynamic combination of EV chargers and single-phase induction motors Implement an LVRT scheme to inject reactive power into the grid and regulate EV charging rate Handle faults and prevent dynamic voltage instability. When the voltage becomes stable, EV loads can be charged at full rate after about 0.6 seconds
almeida2015electric 2015 Combination of inertial emulation and droop control Primary frequency control technique with EVs Safe integration of intermittent renewable energy sources It is verified that EV participation in frequency control in an isolated test system reduced the frequency oscillation band of the system
pandit2024frequency 2024 MM-SFR model Propose a non-linear optimization framework incorporating constraints of EV aggregator, frequency security, converter voltage, security, and LVRT constraints Develop a framework for quantifying EVs’ contribution to providing frequency support Computation time of around 3 s/step
mckillop2023impact 2023 Fault condition test Test EVs physically under various network fault conditions using a grid simulator supply interface Characterize fault ride-through (FRT) performances The aggregation of EV fault-responses yields a resultant successive over-frequency and under-frequency of 50.78 and 49.22 Hz respectively
katic2019impact 2019 V2G mode according to the LVRT requirements Test the possibility of voltage improvement during voltage dips in V2G mode Provide grid support during faults through V2G mode 65.7% higher voltages can be achieved during the dip with LVRT supported by EV
yadav2023low 2023 LVRT Applications of LVRT Boost the resilience of the power network against extreme events With the inclusion of the fault impedance, the grid voltage does not collapse to zero
falahi2013potential 2013 V2G services Pair up a photovoltaic source with an EV charger through a single-phase bidirectional charger topology Study the potential of EVs to help PV sources during LVRT The charger can keep the voltage at the nominal value
dietmannsberger2017simultaneous 2017 A three-phase inverter model Inverters with both LVRT capability and anti-islanding protection simultaneously Solve the conflict between LVRT capability and anti-islanding detection requirements The inverter can behave correctly in all necessary cases, even unbalanced conditions
@@ 3 Management methods of regulating impacts from EV charging
Due to the characteristics of EVs, their batteries can be regarded both as power loads and as distributed energy storage units. By leveraging the storage capability of EV batteries, EVs can provide energy storage, power supply, and ancillary services to the distribution system when parked. With the modernization of power systems, the integration of power and communication infrastructures within the smart grid enables bidirectional energy exchanges and information flows between EVs and the grid, thereby supporting various services that enhance the reliability and sustainability of power systems. This section summarizes the management methods to mitigate the negative impacts in Section 2, including smart charging, charging environment management, energy coordination, battery management, and ancillary services.
@@ 3.1 Smart charging
In response to the growing number of EVs and their impacts on infrastructure, it is necessary to implement intelligent and coordinated charging management methods engel2022hierarchical . Smart charging refers to the intelligent scheduling of EV charging by leveraging data and communication technologies to reduce the adverse impacts of uncontrolled charging, considering grid conditions, electricity prices, and user travel needs dahiwale2024comprehensive . Many studies have focused on optimizing charging management through smart charging approaches, typically integrated with an energy management system (EMS), which is designed to determine optimal charging schemes and regulation strategies that lead to positive effects on the distribution network. Figure 2 presents the negative impacts from EV charging on net loads, and the positive impacts from smart charging. Through establishing models for controllable loads solanki2015including ; hafez2016integrating ; mosaddegh2017optimal , these models are integrated into the distribution system operation framework or EMS solanki2017sustainable ; wi2013electric ; van2016energy ; solanki2015including to determine the optimal charging strategies and dispatch decisions wi2013electric ; van2016energy ; engel2022hierarchical ; solanki2015including ; hafez2016integrating ; kiviluoma2011methodology ; zheng2018online . The advantage of this approach is optimizing EV charging management together with other components of power system. Determining the optimal charging strategies and optimal dispatch decisions is typically formulated as an optimization problem considering multiple factors, with the charging management objective encoded as a cost function engel2022hierarchical . Specifically, in solanki2015including , an EMS framework is proposed to determine optimal scheduling decisions considering dispatchable generators, energy storage systems, and peak demand for controllable loads. zheng2018online transforms the optimal charging problem to an optimal power flow problem to minimize the total system energy cost, then utilizes a modified convex relaxation technique to obtain the globally optimal solution. The key issue in establishing EV load models is reducing the uncertainties caused by EV charging behaviors. Due to the complexity of controllable loads and the limited data, it is difficult to model these loads using basic physical laws mosaddegh2017optimal . Common methods include probabilistic approaches papadopoulos2010predicting ; ramadhani2020review , stochastic optimization van2016energy ; liao2015dispatch , and evolutionary algorithms, such as neural networks hafez2016integrating ; mosaddegh2017optimal , particle swarm optimization regulski2014estimation , or genetic algorithms ju1996nonlinear . Moreover, based on these technical foundations, many studies have focused on developing efficient and accurate real-time management models to address the impact of EV integration on the real-time operation of distribution networks. A detailed introduction to this topic is provided in Section 4.
@@
Figure 2: (a) The effects of EV charging on the electricity load profile on a day in residential distribution grids. (b) The positive impacts of managing EV charging on net loads through PV coordination and smart charging methods. Figures are developed with data from berg2022data .
@@ 3.2 Charging environment management
In addition to charging stations, common EV charging environments include workplaces and households, where rooftop photovoltaic (PV) systems are typically integrated with energy storage systems (ESS). Regarding workplace conditions, the net-zero energy building (NZEB) is a popular topic, which aims to ensure that on-site electricity generation can fully meet the building’s total electricity demand. While the concept of building energy management system (BEMS) is traditionally applied to control heating, ventilation and air conditioning (HVAC) systems and determine operating schedules in order to reduce energy consumption, NZEB requires a more integrated approach where different types of energy sources are interconnected within the building and coupled with the power grid. Typically, this issue begins with modeling energy-efficient building by energy optimization analysis, solar energy and EV batteries are further integrated into the building energy system, followed by optimization analysis for the best design alternative. Three main aspects are mainly considered: total site energy consumption, capital cost, and comfort level alirezaei2016getting . Identifying the optimal design can be challenging due to conflicting objectives, thus multi-objective optimization methods are often required. For example, alirezaei2016getting investigates the issue based on a system of PV panels, EVs, the main battery, and the power grid. Among nearly 1,990 setpoints, 6 points are selected as optimal alternatives. EVs are modeled and incorporated into the energy system as mobile batteries during non-working hours. Comparative simulations demonstrate that this system could reduce grid electricity demand by up to 68%, and lower electricity bills by 62%. engel2022hierarchical proposes a hierarchical economic model predictive control scheme for EV charging management, considering the objectives of monetary costs, building temperature comfort, EV charge satisfaction, and battery degradation. van2016energy designs a workplace energy management system with photovoltaic generation prediction and power flow optimization between PV systems, grids, and battery electric vehicles. lakshminarayanan2018real proposes a real-time optimal EM controller for V2G integration to provide an optimal schedule for the operation of the workplace microgrid system.
The issue of EV charging in households is closely related to the home energy management system (HEMS), which is necessary for residential electricity consumers to participate in DR programs actively ghazvini2017demand . This issue is typically formulated as an optimization problem, often modeled using a Markov decision process. The objective is to minimize the occupant’s utility function while considering constraints at multiple levels, such as occupant, residential home and distribution grid levels liu2021pv . However, due to the difficulty in accurately quantifying occupant behavior, the effectiveness of proposed strategies is highly dependent on assumed scenarios. For example, seal2023centralized applies a centralized model predictive control (MPC) strategy with zone-based control to manage a heating system comprising a heat pump with multi-split fan units and electric baseboards while integrating PV generation and EV energy storage. The responsiveness of the MPC is evaluated in a vehicle-to-home (V2H) case study where EV arrival time is only notified a few minutes before arrival. garifi2019stochastic proposes a chance-constrained MPC algorithm to manage controllable resources, including PV panels, home batteries, EVs, and HVAC systems, with the goal to ensure indoor thermal comfort despite uncertainties in temperature and solar irradiance forecasts. alirezaei2016getting compares three scenarios differing in EV energy operations and PV placements, and finds the grid electricity consumption can be reduced up to 45% and 77%. irfan2024novel develops a base case to analyze grid power-sharing based on a grid-assisted bidirectional PV–EV system using the system advisor model and conducts tests in Sydney households. liu2021pv develops a stochastic adaptive dynamic programming model to optimize HVAC setpoints, clothing behavior, and EV energy scheduling, accounting for uncertainties in outdoor temperature, PV generation, and EV’s state of charge (SOC).
Some studies investigate the optimal energy flow that motivates these scenarios, focusing on balancing thermal comfort, electricity cost minimization, and the integration of distributed energy resources. For instance, seal2023centralized develops an MPC-based HEMS to manage zone-based thermal comfort along with optimizing the energy flow among the components of the home energy network. Specifically, the MPC optimizes heating system inputs to minimize energy cost, including the part load ratio and the percentage of the rated baseboard heating input. A multistep MPC feedback strategy is employed with a simulation time step of 3.75 min, a prediction horizon of 8 h, and a control horizon of 15 min. The MPC also achieves approximately 8% reduction in the energy cost compared with the base-case scenario. irfan2024novel proposes a data-driven HEMS model based on the proximal policy optimization algorithm to optimize policy formulation in sequential decision-making tasks. It prioritizes the lowest-cost energy sources and leverages V2H and V2G functionalities to minimize monthly electricity costs. In van2016energy , an EMS that combines an autoregressive integrated moving average model for PV forecasting and a mixed-integer linear programming framework for power allocation is employed to optimize power flows among PV, EV, and the grid. This approach reduces EV charging costs by 118.44% with one charging point and 427.45% with two charging points compared with uncontrolled charging. liu2021pv introduces HEMS based on adaptive dynamic programming. A model predictive control framework is further integrated to continually update optimal appliance scheduling decisions under time-of-use tariff, achieving a 68.5% reduction in energy costs compared with conventional scheduling strategies.